OpenStudy (anonymous):

What is an asymptote of the hyperbola given by the equation [(y^2)/(1)]-[(x^2)/(81)]=1

OpenStudy (loser66):

hey hey cute cat, you have formula for this, just apply

OpenStudy (anonymous):

Sorry, I'm a dog not a cat :3

OpenStudy (loser66):

oh yeah.... cute dog!!! look like a cat

OpenStudy (loser66):

as I told , you have formula, apply.

OpenStudy (anonymous):

cept i don't what is the formula?

OpenStudy (loser66):

with a hyperbola \[\frac{y^2}{a} -\frac{x^2}{b}=1\] the asymptotes y = +- (a/b)x your a =1 your b =9

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

@loser66

OpenStudy (loser66):

what you mean by "then what?" then apply to that formula to get y = 1/9 x and y = -1/9 x . they are 2 asymptotes of your hyperbola.

OpenStudy (anonymous):

y = 1/9x y = 9x y = 1/81x y = 81x

OpenStudy (loser66):



OpenStudy (loser66):

got what I mean?

OpenStudy (anonymous):

no

OpenStudy (loser66):

@Mertsj

OpenStudy (loser66):

so, wait for Mertsj. me done

OpenStudy (anonymous):

just forget it